OpenStudy (anonymous):

A model scale is 1 cm = 8 m. If the actual object is 40 meters, the model is _____ cm long.

OpenStudy (nirmalnema):

5 cm

OpenStudy (anonymous):

@nirmalnema have them solve it out dont give the answer directly

OpenStudy (nirmalnema):

1cm = 8m 1m = 1/8 cm 40m = 1/8 * 40 =5

OpenStudy (texaschic101):

1/8 = x/40 --- 1 cm to 8 m = x cm to 40 m cross multiply 8x = 40 -- divide both sides by 8 x = 5 cm

OpenStudy (anonymous):

yes what they said... @texaschic101 had the better example because it uses proportions so it is 5